Emily is 6 times as old as her daughter. Next year she will be 5 times her daughter's age. find the present age of each and check the answer
Answer by mananth(16946)

daughter = x
emily = 6x
after 1 year
daughter = (x+1)
Emily = (6x+1)
6x+1=5(x+1)
6x+1=5x+5
6x-5x=5-1
x=4 daughter's age
Emily = 6x= 24 years